### v1

# Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

Abstract
Masked Image Modeling (MIM) has emerged as a pivotal approach for developing foundational visual models in the field of remote sensing (RS). However, current RS datasets are limited in volume and diversity, which significantly constrains the capacity of MIM methods to learn generalizable representations. In this study, we introduce RS-4M, a large-scale dataset designed to enable highly efficient MIM training on RS images. RS-4M comprises 4 million optical images encompassing abundant and fine-grained RS visual tasks, including object-level detection and pixel-level segmentation. Compared to natural images, RS images often contain massive redundant background pixels, which limits the training efficiency of the conventional MIM models. To address this, we propose an efficient MIM method, termed SelectiveMAE, which dynamically encodes and reconstructs a subset of patch tokens selected based on their semantic richness. SelectiveMAE roots in a progressive semantic token selection module, which evolves from reconstructing semantically analogical tokens to encoding complementary semantic dependencies. This approach transforms conventional MIM training into a progressive feature learning process, enabling SelectiveMAE to efficiently learn robust representations of RS images. Extensive experiments show that SelectiveMAE significantly boosts training efficiency by 2.2-2.7 times and enhances the classification, detection, and segmentation performance of the baseline MIM model. The dataset, source code, and trained models will be released at RS-4M.
1 Introduction
Over the past decade, advancements in remote sensing (RS) technology and data acquisition have significantly improved applications in ecosystem monitoring [1], natural disaster management [2], among others [3, 4]. These applications rely on essential capabilities such as scene classification [5, 6], object detection [7], change detection [8], and semantic segmentation [9]. However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
Thanks to significant advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [10, 11], the pre-training of visual foundation models has seen remarkable improvements [12, 13, 14, 15, 16, 17]. Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [18]. However, two challenges persist in the development of RSFMs. (i) Compared to the ImageNet-21k [19] dataset, previous RS datasets [20, 21, 22, 23] contain significantly fewer samples (approximately 1 million vs. 14 million), which impedes the sufficient MIM training of large backbones. (ii) These datasets, which primarily focus on global scene semantics [20, 21, 22], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks. This limitation restricts the generalization of the learned representations.
To address these challenges, we introduce a large-scale RS dataset called RS-4M. This dataset, comprising 4 million optical images, is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. RS-4M significantly exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger. Moreover, RS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 1 left).
Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the significant computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [39] backbone on 8 Nvidia A100 GPUs [18]. In natural scene analysis, this issue has led to numerous studies [40, 41, 42, 43, 44, 45] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [40, 41]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], significantly speeding up feature extraction.
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [18, 28]. This raises two key questions about how to efficiently conduct MIM training in the RS field: (i) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? (ii) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
We propose an efficient MIM method called SelectiveMAE, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens ($e.g.,\leq$50%) with higher HOG values for feature encoding ($e.g.,\leq$15%) and pixel reconstruction ($e.g.,\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing (Fig. 3). To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically leverages semantically relevant patch tokens throughout the entire MIM training phase. At the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary approach allows SelectiveMAE to efficiently and progressively learn robust representations of RS images, thereby accelerating MIM convergence. Our experiments reveal that only 40% of RS image patches are sufficient to train a comparable MIM model, offering new insights into conducting MIM training on RS images.
The contributions of this paper are summarized as follows. (1) We introduce the RS-4M dataset, a large-scale optical remote sensing dataset for unsupervised learning methods to date, distinguished by its diverse scene details. (2) We propose SelectiveMAE, a novel and efficient MIM method tailored for remote sensing images. This method incorporates a new PSTS module, which significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
Experiment results suggest that compared to the baseline method, SelectiveMAE not only achieves 2.22.7×2.2-2.7$\times$2.2 - 2.7 × speedup in pre-training but also provides performance improvement of 5.6% mAP and 1.4% mF1 when applied to downstream object detection and segmentation tasks, respectively.
2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [23, 46], object detection [47, 48, 49], and segmentation [50, 51, 52]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works, like SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches, combine various sensor data to create extensive datasets, while many other datasets focus solely on optical RS images. MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times. These datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, we introduce the RS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
Remote Sensing Foundation Models. remains unlabeled and thus inaccessible for supervised learning [54]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Although there are some contrastive self-supervised methods [24, 21, 22, 55], recent advancements have primarily centered around generative self-supervised methods. For example, RVSA [18] leverages the vision transformer backbone pre-trained by the vanilla MAE [10] method and introduces computationally efficient rotated varied-size window attention to replace the original attention during fine-tuning. In addition, many studies are focusing on improving generative self-supervised algorithms by leveraging general image knowledge [25, 26, 28], scaling up parameter sizes [29], integrating spatio-temporal data [30, 31, 32], handling multi-sensor data [56, 57, 58, 59, 60], and employing multi-scale concepts [33, 35, 36]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In this paper, we propose a novel acceleration method that significantly speeds up training, enhancing practicality in the RS domain.
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [61], MIM has been developed for visual pre-training [62, 11, 10, 63, 64]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [65, 66, 67, 68, 69], masking strategies [70, 71], and reconstruction methods [72, 73, 74, 75]. For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or ues a novel Difficulty-Flatten Loss [43]. Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. In this paper, we introduce a novel PSTS module to address this issue and accelerate the training process.
3 Dataset
3.1 Data collection and organization
Figure 1: Left: The organization of data sources in RS-4M. Right: The comparison between RS-4M and other datasets.
Recent advancements in self-supervised pre-training RSFMs are hindered by the smaller scale and diversity of existing RS datasets compared to natural scene datasets. To address this issue, we propose the RS-4M dataset, a continuously updated large-scale RS dataset following the principles of Diversity, Richness, and Scalability (DiRS) [20]. (1) We collected and organized datasets from the past decade, focusing on mainstream RS image interpretation tasks, including scene classification, object detection, semantic segmentation, and change detection. These datasets, detailed in Fig. 1, encompass around 100 sub-datasets with a total of 13.5 million images. (2) We excluded multispectral and SAR data, keeping only optical images in this study. Future updates will incorporate multispectral and SAR data to support multi-modal self-supervised pre-training. (3) Unlike datasets like MillionAID [20] designed for scene classification, our RS dataset is designed to facilitate downstream detection and segmentation tasks. To address large-scale image in segmentation datasets, we randomly cropped high-resolution images into smaller slices. (4) Images with pixels below 64 or above 1024 were removed to facilitate self-supervised pre-training. (5) We combined the remaining images and eliminated duplicates using a two-phase process: a coarse phase with perceptual hashing[76] and a refined phase with manual review. With the Hash value to select and the checking of manual review, only highly similar images are excluded. This resulted in the RS-4M dataset containing about 4 million high-quality RS optical images, which is four times larger than previous representative RS datasets, as shown in Fig. 1 right.
Table 1: Comparison between RS-4M and MillionAID: ViT Base models [39] are pre-trained on these datasets using MAE [10] and fine-tuned on the RESISC-45 dataset [77].
Dataset Images Number Epoch Top-1 Accuracy
MillionAID [20] 1 Million 800 89.20
RS-4M 2 Million 400 91.80
RS-4M 3 Million 267 92.24
RS-4M 4 Million 200 92.38
3.2 Preliminary Evaluation of RS-4M Dataset
RS-4M offers a significantly larger and more diverse image set compared to previous datasets. To evaluate its effectiveness, we pre-train a ViT base model [39] using the vanilla MAE method [10]. For comparison, we use the MillionAID dataset, maintaining an equal number of data points during training: 800 epochs for MillionAID’s 1 million images and 200 epochs for our RS-4M dataset. In addition, we create new datasets by sampling different numbers of images from RS-4M for further comparison. After pre-training, we fine-tune the models on the downstream RESISC-45 dataset [77] and assess the Top-1 classification accuracy. The results in Table 1 demonstrate that the RS-4M dataset outperforms MillionAID across various sample sizes, with the highest accuracy of 92.38% achieved using the entire dataset. Even with fewer images, RS-4M consistently delivers superior performance, highlighting its greater diversity and effectiveness. These findings underscore the advantage of using larger and more varied datasets for unsupervised pre-training.
4 Method
4.1 Masked Autoencoders Preliminaries
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy. In MAE, the encoder processes only 25% of the patches, significantly reducing computational load by avoiding processing the entire image. Building on MAE, we aim to leverage the redundancy in RS images to accelerate training. Specifically, we address two questions: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
Figure 2: Overview of SelectiveMAE. The vanilla version of SelectiveMAE.
4.2 Partial Reconstruction
Table 2: Results of different partial reconstruction strategies. Reconstruction Ratio=25%. Top-1 classification accuracy on RESISC-45 [77] is reported.
Method Selection Data Throu. Acc.
/Min.
Partial Reco. False 17k 89.2
Partial Reco. Random 25k 88.8
Partial Reco. ATS [78] 20k 85.7
Partial Reco. HOG 25k 89.2
For question 1, previous research [45] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches. In fact, reconstructing just 50% or even 25% of the patches can achieve similar performance and speed up training. However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones. As shown in Table 2, using only a random subset for reconstruction degrades performance.
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 4.1. The remaining patches serve as reconstruction targets for the decoder. Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
[…]
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [45]. Experimental results in Table 2 show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
4.3 Progressive Semantic Token Selection
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 4.2. However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig. 3. To illustrate, we continued MAE’s training despite these gradient explosions. The figure shows that using only 40% (=15%+25%) of the patches for encoding and reconstruction frequently caused unstable training due to gradient explosions. We speculate that, although RS images have background redundancy, using such a small portion of patches significantly challenges MAE training.
Figure 3: Left: Baseline training loss. Using only 40% of masked patches for reconstruction often leads to gradient explosions. Right: Training loss using PSTS.
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Some studies suggest that machine learning algorithms often benefit from starting with simpler tasks and gradually increasing the difficulty, a strategy known as curriculum learning [79, 80, 81]. This approach involves progressively introducing more challenging examples during training. While curriculum learning typically focuses on whole image samples, self-supervised training, which masks large portions of image pixels, emphasizes patch-level selection. Therefore, in self-supervised training, it’s essential to distinguish between “easy" and “hard" patches.
To this end, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection. In this module, depicted in Fig. 2, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. We define the initial set of selected token as:
[…]
We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set. This simple yet effective strategy ensures that semantically rich tokens are selected.
Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens. This process is outlined in Algorithm 1. Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values. Recalling our partial reconstruction approach in Section 4.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets. Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process. Conversely, if a token’s feature distribution significantly deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn.
Specifically, we select tokens from $S^{U}$ based on $S^{I}$. First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer. We use Cosine Distance to measure the distance between the tokens in these two sets:
$\mathcal{D}(S^{U},S^{I})=\mathbf{1}-cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-1.07639pt\left|\mathbb{S^{U}}\right|\kern-1.07639pt\right|\cdot\left|\kern-1.07639pt\left|\mathbb{S^{I}}\right|\kern-1.07639pt\right|},$ (3)
where $\mathbf{1}$ is an all-one matrix. $\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$. Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
[…]
where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs. Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
[…]
${S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$. The operations in Eq. (5) and Eq. (6) are performed in each training epoch. This process is summarized in Algorithm 1.
Algorithm 1 Progressive Semantic Token Selection
1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
3:for $t\leftarrow 1$ to $T$ do
4:     Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
5:     $s\leftarrow\frac{1}{2}(1-m)$
6:     Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
7:     Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
[…]
9:     Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
10:end for
5 Experiment
In this section, we first discuss the pre-training process using the RS-4M dataset, comparing the training speed and performance of our proposed SelectiveMAE with existing MAE acceleration methods. We then demonstrate SelectiveMAE’s superior performance across various downstream tasks compared to different RSFMs. Finally, we explore design choices such as the reconstruction ratio, mask ratio, and decoder depth through extensive ablation experiments, detailed in the Appendix.
5.1 Pre-training
Pre-training Setup. Our pre-training experiment setup largely follows MAE [10]. Unlike other MIM methods in the RS field, such as SatMAE [30] and ScaleMAE [33], we train both the commonly used ViT-B and the larger ViT-L backbones [39]. We adjust the input image to a size of $224\times 224$, with a patch size of 16, using the AdamW optimizer and a cosine learning rate scheduler. The initial learning rate is set to 1.5e-4, batch size to 1024, masking ratio to 85%, and reconstruction ratio to 25%. For patch selection during the 800 epochs, the adjustment thresholds are set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
Table 3: Top-1 classification accuracy on RESISC-45 [77]. “-” indicates unavailable accuracy data. Training throughput was measured using 8 NVIDIA 40G A100 GPUs.
Model Publication Backbone Params Data Throughput/Min. Acccuracy
MAE [10] CVPR’22 ViT-B 86M 17k 93.45
CrossMAE [45] Arxiv’24 ViT-B 86M 22k 92.82
SatMAE [30] NIPS’22 ViT-B 86M 15k -
ScaleMAE [33] ICCV’23 ViT-B 86M 16k -
SelectiveMAE - ViT-B 86M 37k ($\times2$) 94.58 ($+1.13$)
MAE [10] CVPR’22 ViT-L 307M 15k 93.80
SatMAE [30] NIPS’22 ViT-L 307M 13k 94.10
ScaleMAE [33] ICCV’23 ViT-L 307M 13k 95.04
SelectiveMAE - ViT-L 307M 41k (2.7×2.7$\times$2.7 ×) 95.77 ($+1.97$)
Pre-training Results.
We pre-trained ViT-B and ViT-L on the RS-4M dataset, evaluating the image data throughput per minute for SelectiveMAE and the baseline (MAE) [10]. We also assessed the training speed of the general domain acceleration method CrossMAE [45], as well as the remote sensing methods SatMAE [30] and ScaleMAE [33]. Table 3 illustrates our method’s more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline. It also outperforms the general domain acceleration method CrossMAE significantly. Additionally, we reported the top-1 classification accuracy on the same dataset, with both MAE and CrossMAE trained on MillionAID [20]. Table 3 clearly demonstrates our method’s superior speed and performance.
The significant acceleration effect of SelectiveMAE stems from utilizing only 40% of the image patches (15% for the encoder and 25% for the decoder). Despite this reduced utilization, it achieves superior performance compared to other acceleration methods [45] due to the proposed novel token masking and reconstruction strategy, tailored for the remote sensing images. This unique approach ensures that SelectiveMAE delivers both speed and high performance, as validated in Table 4.
5.2 Fine-tuning on Downstream Tasks
We further evaluated SelectiveMAE across three key downstream tasks: scene classification, object detection, and semantic segmentation. In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[82], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models. These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning based approaches [83]. Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE). To ensure fairness, all experiments followed to consistent fine-tuning settings. Supplementary materials and code provide comprehensive experimental details.
Table 4: Results for scene classification, object detection, and semantic segmentation. “TR" represents the ratio of training data to the entire dataset.
Model Publication Backbone Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
SeCo [22] ICCV’21 ResNet-50 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [24] ICCV’21 ResNet-50 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [28] JSTARS’23 ResNet-50 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [21] CVPR’23 ResNet-50 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [30] NIPS’22 ViT-L 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [33] ICCV’23 ViT-L 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [82] GRSM’23 ViT-S 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [37] TGRS’22 Swin-B 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [83] ICCV’23 Swin-B 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [25] ICCV’23 Swin-B 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [18] TGRS’23 ViT-B+RVSA 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
SelectiveMAE - ViT-B 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
SelectiveMAE - ViT-L 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
Scene Classification. We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders. We leverage two scene classification datasets: AID [84] and RESISC-45 [77]. Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code. Evaluation is based on overall accuracy (OA). Table 4 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline.
Besides, it outperforms RVSA and others when scaled to ViT-L. This demonstrates that, with our SelectiveMAE training method on the RS-4M dataset, the model can learn strong feature representations and scale efficiently.
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [85] and its improved variant DIOR-R for oriented object detection [86], both comprising RGB images. Following the methodologies of previous works [18, 37], we maintained consistent experimental setups, employing Faster-RCNN [87] and Oriented-RCNN [88] as detectors for each dataset. Detailed experimental configurations are available in the accompanying code. The results are summarized in Table 4. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37], SatLas [83], and GFM [25]. This is particularly noteworthy as hierarchical transformers generally excel at generating multiscale features essential for object detection. Moreover, our SelectiveMAE with a ViT-B backbone outperforms SatMAE [30] and ScaleMAE [33], which use a larger ViT-L backbone, and matches the advanced RVSA [18, 34] that employs specialized window attention. When using the larger ViT-L backbone, SelectiveMAE shows significantly enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [51] and SpaceNetv1 [52]. Our implementation closely follows [18], utilizing UperNet [89] as the segmentation framework, with mean Intersection over Union (mIOU) as the evaluation metric. Table 4 demonstrates the clear superiority of SelectiveMAE over its competitors in pixel-level semantic segmentation. The selection and reconstruction strategies employed by SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
6 Conclusion
In this paper, we introduce RS-4M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous datasets in the RS field, RS-4M offers a significantly larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on RS-4M highlights its advantages in these tasks. To address the computational burden of MIM training on large-scale RS datasets, we propose SelectiveMAE, an efficient MIM method that dynamically encodes and reconstructs patch tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that only 40% of RS image patches are needed to train a comparable MIM model. Experiments show that SelectiveMAE not only achieves over 2×2$\times$2 × speedup in pre-training but also enhances performance in various downstream tasks.
Limitations and Discussion. Although the RS-4M dataset surpasses existing RS datasets by more than four times, its size can be further increased by collecting and filtering additional RS data. Additionally, its diversity can be enhanced by incorporating data from various types of sensors.
    A: The dataset was created by:
    - Fengxiang Wang (National University of Defense Technology),
    - Hongzhen Wang (Tsinghua University),
    - Di Wang (Wuhan University),
    - Zonghao Guo (University of Chinese Academic of Sciences),
    - Zhenyu Zhong (Nankai University),
    - Long Lan (National University of Defense Technology),
    - Jing Zhang (The University of Sydney),
    - Zhiyuan Liu (Tsinghua University),
    - Maosong Sun (Tsinghua University).
  3. 3.
    “Who funded the creation of the dataset?”
    A: The dataset creation was funded by the affiliations of the authors involved in this work.
7.5.2 Composition
Most of the questions in this section are intended to provide dataset consumers with the information they need to make informed decisions about using the dataset for their chosen tasks. Some of the questions are designed to elicit information about compliance with the EU’s General Data Protection Regulation (GDPR) or comparable regulations in other jurisdictions. Questions that apply only to datasets that relate to people are grouped together at the end of the section. We recommend taking a broad interpretation of whether a dataset relates to people. For example, any dataset containing text that was written by people relates to people.
  1. 1.
    “What do the instances that comprise our datasets represent (e.g., documents, photos, people, countries)?”
    A: The dataset primarily comprises visible light remote sensing images captured by satellites. All datasets utilized in RS-4M are publicly accessible and nonprofit.
  2. 2.
    “How many instances are there in total (of each type, if appropriate)?”
    A: RS-4M contains 4 million remote sensing image instances captured by satellites.
  3. 3.
    “Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?”
    A: Yes, our dataset contains all possible instances that have been collected so far.
  4. 4.
    “Is there a label or target associated with each instance?”
    A: No, our dataset is intended for self-supervised learning. Therefore, each instance is an individual remote sensing image and does not contain annotations.
  5. 5.
    “Is any information missing from individual instances?”
    A: No.
  6. 6.
    “Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?”
    A: Yes, the relationship between individual instances is explicit.
  7. 7.
    “Are there recommended data splits (e.g., training, development/validation, testing)?”
    A: Yes, the entire dataset is intended for self-supervised methods, and we recommend using the whole dataset for self-supervised learning research.
  8. 8.
    “Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?”
    A: Yes, our dataset relies on many publicly available remote sensing datasets, which we have detailed in the main text.
  9. 9.
    “Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)?”
    A: No, all data are clearly licensed.
  10. 10.
    “Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?”
    A: No.
7.5.3 Collection Process
In addition to the goals outlined in the previous section, the questions in this section are designed to elicit information that may help researchers and practitioners create alternative datasets with similar characteristics. Again, questions that apply only to datasets that relate to people are grouped together at the end of the section.
  1. 1.
    “How was the data associated with each instance acquired?”
    A: Please refer to the details listed in the main text Sec. 3.
  2. 2.
    “What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?”
    A: Please refer to the details listed in the main text Sec. 3.
  3. 3.
    “If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?”
    A: Please refer to the details listed in the main text Sec. 3.
7.5.4 Preprocessing, Cleaning, and Labeling
The questions in this section are intended to provide dataset consumers with the information they need to determine whether the “raw” data has been processed in ways that are compatible with their chosen tasks. For example, text that has been converted into a “bag-of-words" is not suitable for tasks involving word order.
  1. 1.
    “Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?”
    A: Yes, we preprocessed and cleaned data in our dataset.
  2. 2.
    “Was the ‘raw’ data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?”
    A: Yes, raw data is accessible.
  3. 3.
    “Is the software that was used to preprocess/clean/label the data available?”
    A: Yes, the necessary software used to preprocess and clean the data is publicly available.
7.5.5 Uses
The questions in this section are intended to encourage dataset creators to reflect on tasks for which the dataset should and should not be used. By explicitly highlighting these tasks, dataset creators can help dataset consumers make informed decisions, thereby avoiding potential risks or harms.
  1. 1.
    “Has the dataset been used for any tasks already?”
    A: No.
  2. 2.
    “Is there a repository that links to any or all papers or systems that use the dataset?”
    A: Yes, we provide such links in our GitHub repository.
  3. 3.
    “What (other) tasks could the dataset be used for?”
    A: The dataset could be used for training the remote sensing foundation models with the self-supervised learning method.
  4. 4.
    “Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?”
    A: N/A.
  5. 5.
    “Are there tasks for which the dataset should not be used?”
    A: N/A.
7.5.6 Distribution
Dataset creators should provide answers to these questions prior to distributing the dataset either internally within the entity on behalf of which the dataset was created or externally to third parties.
  1. 1.
    “Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?”
    A: No.
  2. 2.
    “How will the dataset be distributed (e.g., tarball on website, API, GitHub)?”
    A: Very likely to be distributed by website, API, and GitHub repository.
  3. 3.
    “When will the dataset be distributed?”
    A: The datasets are publicly accessible.
  4. 4.
    “Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?”
    A: Yes, the dataset is under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License.
  5. 5.
    “Have any third parties imposed IP-based or other restrictions on the data associated with the instances?”
    A: No.
  6. 6.
    “Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?”
    A: No.
7.5.7 Maintenance
As with the questions in the previous section, dataset creators should provide answers to these questions prior to distributing the dataset. The questions in this section are intended to encourage dataset creators to plan for dataset maintenance and communicate this plan to dataset consumers.
  1. 1.
    “Who will be supporting/hosting/maintaining the dataset?”
    A: The authors of this work serve to support, host, and maintain the datasets.
  2. 2.
    “How can the owner/curator/manager of the dataset be contacted (e.g., email address)?”
    A: The curators can be contacted via the email addresses listed on our webpage [^4] https://github.com/Fengxiang23/SelectiveMAE..
  3. 3.
    “Is there an erratum?”
    A: There is no explicit erratum; updates and known errors will be specified in future versions.
  4. 4.
    “Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?”
    A: Yes, for the current version. Future updates (if any) will be posted on the dataset website.
  5. 5.
    “Will older versions of the dataset continue to be supported/hosted/maintained?”
    A: Yes. This is the first version of the release; future updates will be posted and older versions will be replaced.
  6. 6.
    “If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?”
    A: Yes, we provide detailed instructions for future extensions.
7.6 Limitation and Potential Societal Impact
In this section, we elaborate on the limitations and potential societal impact of this work.
7.6.1 Potential Limitations
While RS-4M provides a comprehensive benchmark for training the remote sensing foundation models with self-supervised learning methods, there are several limitations to consider:
  • Scope of Sensors: Although our benchmark includes 4 million visible light remote sensing images, it may not cover all possible real-world scenarios. There could be additional sensor data, like multispectral data that were not included in this study, potentially limiting the generalizability of our findings.
  • Model and Dataset Diversity: While our dataset is primarily focused on adapting to downstream tasks like detection and segmentation, there is undeniably less data available for these tasks compared to general scene classification. In the future, we should collect more data that is better suited for various downstream tasks.
  • Computation and Resource Requirements: Pre-training on the extensive RS-4M dataset, comprising 4 million images, demands substantial computational resources, despite our introduction of the efficient SelectiveMAE method. This may limit access to the benchmark for research groups lacking ample computational power.
7.6.2 Potential Negative Societal Impact
While the development of remote sensing foundation models with self-supervised learning methods has the potential to significantly advance remote sensing downstream tasks, there are potential negative societal impacts that must be considered:
  • Safety Risks: Our benchmark aims to improve remote sensing foundation models, but relying too heavily on these models may breed overconfidence in autonomous systems. It’s crucial to deploy these systems with adequate safety measures and human oversight to uphold public safety.
  • Environmental Impact: Training and evaluating remote sensing foundation models demand substantial computational resources, leading to a notable environmental footprint. Encouraging the adoption of energy-efficient algorithms (i.e., the proposed SelectiveMAE) and sustainable computing practices is crucial to reduce the environmental impact of this research.
  • Bias and Fairness: The performance of foundational remote sensing models can vary across different environments and conditions, potentially introducing biases in downstream tasks within the remote sensing domain. It is essential to train and evaluate these models on diverse datasets to ensure fairness and avoid discriminatory outcomes.