### v3

# OpticalRS-4M: Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-4M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-4M comprises 4 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models. The dataset, source code, and trained models will be released at OpticalRS-4M.
1 Introduction
Over the past decade, advancements in remote sensing (RS) technology and data acquisition have improved applications in ecosystem monitoring [1], natural disaster management [2], among others [3, 4]. These applications rely on essential capabilities such as scene classification [5, 6], object detection [7], change detection [8], and semantic segmentation [9]. However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
Thanks to advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [10, 11], the pre-training of visual foundation models has seen remarkable improvements [12, 13, 14, 15, 16, 17]. Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [18]. However, two challenges persist in the development of RSFMs. (i) Compared to the ImageNet-21k [19] dataset, previous RS datasets [20, 21, 22, 23] contain fewer samples (approximately 1 million vs. 14 million), which impedes the sufficient MIM training of large backbones. (ii) These datasets, which primarily focus on global scene semantics [20, 21, 22], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks. This limitation restricts the generalization of the learned representations.
To address these challenges, we propose a new pipeline to collect, create, and efficiently process a large optical RS dataset. Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [20]. Nevertheless, issues such as inconsistent data sources, excessively large image sizes, and redundant pixels still exist. Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality. Correspondingly, we obtain a large-scale RS dataset called OpticalRS-4M comprising 4 million optical images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. OpticalRS-4M exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger. Moreover, OpticalRS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 1 left).
Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [39] backbone on 8 Nvidia A100 GPUs [18]. This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-4M. In natural scene analysis, this issue has led to numerous studies [40, 41, 42, 43, 44, 45] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [40, 41]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], speeding up feature extraction.
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [18, 28]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens ($e.g.,\leq$50%) with higher HOG values for feature encoding ($e.g.,\leq$15%) and pixel reconstruction ($e.g.,\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence. Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
The contributions of this paper are summarized as follows:
1) We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
2) Using this pipeline, we create the OpticalRS-4M dataset, which is a large-scale RS dataset comprising 4 million optical images with diversified coverage scenarios.
3) We introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training. It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
Experiment results suggest the effectiveness and scalability of the proposed pipeline. OpticalRS-4M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [10].
2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [23, 46], object detection [47, 48, 49], and segmentation [50, 51, 52]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches. SSL4EO-S12 [54] dataset uses publicly available SAR/optical satellite data collected by the European Space Agency’s Sentinel mission. SSL4EO-L [55] consists entirely of multispectral data. SatlasPretrain [56] is constructed from data sources such as Sentinel-2, Sentinel-1, Landsat 8/9, and 0.5 - 2 meter/pixel aerial imagery. While it includes millions of images, the majority are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [57] and RS5M [58], but they primarily focus on multimodal tasks. Currently, some datasets focus solely on optical RS images. MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, with the proposed pipeline, we introduce the OpticalRS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [59], MIM has been developed for visual pre-training [60, 11, 10, 61, 62]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [63, 64, 65, 66, 67], masking strategies [68, 69], and reconstruction methods [70, 71, 72, 73]. For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or use a novel Difficulty-Flatten Loss [43]. Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. Considering these issues, we introduce an efficient MIM method named SelectiveMAE that significantly speeds up pre-training, enhancing the practicality of developing RS foundation models based on large-scale datasets.
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [74]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Although there are some contrastive self-supervised methods [24, 21, 22, 75], recent advancements have primarily centered around generative self-supervised methods. For example, RVSA [18] leverages the vision transformer backbone pre-trained by the vanilla MAE [10] method and introduces computationally efficient rotated varied-size window attention to replace the original attention during fine-tuning. In addition, many studies are focusing on improving generative self-supervised algorithms by leveraging general image knowledge [25, 26, 28], scaling up parameter sizes [29], integrating spatio-temporal data [30, 31, 32], handling multi-sensor data [76, 77, 78, 79, 80], and employing multi-scale concepts [33, 35, 36]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
3 Methodology
The proposed pipeline contains two critical components: dataset generation and efficient pre-training, which will be introduced in detail in the following text.
3.1 Dataset Generation
Figure 1: Left: The organization of data sources in OpticalRS-4M. Right: The comparison between OpticalRS-4M and other datasets.
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this, we first curate a large-scale RS dataset, OpticalRS-4M, through extensive data collection and preprocessing.
3.1.1 Data Collection
We began by reviewing publicly available RS datasets from the past decade. Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [20]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity. Intra-class Diversity: High diversity within classes ensures comprehensive representation of real-world target distributions. To ensure this, we consider the data from different sources of the same scene or target. Moreover, in the RS field, the attributes of objects may vary with geographic location and imaging time, so the change detection data is also incorporated. Inter-class Diversity: To enable self-supervised pre-training to distinguish between categories effectively, we incorporated as many fine-grained categories as possible, especially those with high semantic overlap. This strategy helps the model learn robust feature representations. 2) Richness: Beyond diversity, richness is also essential. We ensure diverse content features and large sample sizes by collecting images under various conditions — weather, seasons, lighting, and so on — introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion. Our data collection process actively includes the sample of the same category under different imaging conditions to ensure comprehensive real-world representation, enhancing the model’s representation and generalization capabilities. 3) Scalability: Given the evolving applications of remote sensing images, scalability is also crucial. We constantly enrich the dataset based on scene and target categories across classification, detection, and segmentation tasks.
In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios. Fig. 1 left illustrates the included RS datasets collected.
3.1.2 Data Preprocessing
Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained. To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
1) We focused exclusively on optical images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training. 2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices. 3) We filtered out images with fewer than 64 or more than 1024 pixels. 4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [81], and then a refined phase involving manual review. This ensured that only highly similar images were excluded. As a result of these steps, we created the OpticalRS-4M dataset, which contains approximately 4 million high-quality RS images — significantly larger than previous representative RS datasets, as illustrated in Fig. 1 right.
In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality results. We’ve screened and processed most small and medium-scale datasets from the past decade in this field. More efforts can be made to extend OpticalRS-4M by including larger datasets like SSL4EO-S12 [54] and SSL4EO-L [55] to enhance dataset capacity.
3.1.3 Dataset Diversity
Refer to captionRefer to caption
Figure 2: Left: Main categories with visual examples from OpticalRS-4M.Right: Full hierarchy of OpticalRS-4M categories.
To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, reviewing all the scenarios and targets included in OpticalRS-4M, as shown in Fig. 2. The analysis reveals that OpticalRS-4M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical remote sensing datasets, OpticalRS-4M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Concert”, “Parade/Protest”, “Religious Activity”, “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and “Traffic Collision”. This expanded information enhances the versatility of OpticalRS-4M for a wider range of downstream tasks.
3.2 Efficient Pre-training
After obtaining a large-scale RS dataset, the next step is to conduct the self-supervised pre-training. However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [10]. To address this issue, our pipeline introduces an efficient MIM method, i.e., SelectiveMAE. In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
3.2.1 Masked Autoencoders Preliminaries
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables the training of large models with reduced computational and memory requirements. 3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy. In MAE, the encoder processes only 25% of the patches, reducing computational load by avoiding processing the entire image. Building on MAE, we aim to leverage the redundancy in RS images to accelerate training. Specifically, we address two questions: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
Figure 3: Overview of SelectiveMAE. The vanilla version of SelectiveMAE.
3.2.2 Partial Reconstruction
For question 1, previous research [45] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches. In fact, reconstructing just 50% or even 25% of the patches can achieve similar performance and speed up training. However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones. Using only a random subset for reconstruction even degrades performance.
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 3.2.1. The remaining patches serve as reconstruction targets for the decoder. Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
[…]
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [45]. Experimental results in the Appendix show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
3.2.3 Progressive Semantic Token Selection
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 3.2.2. However, during training, this setup often led to issues like gradient explosions or loss divergence (see the Appendix). We speculate that, although RS images have background redundancy, using such a small portion of patches challenges MAE training. How can we achieve acceleration at a high masking ratio while ensuring MAE convergence? Curriculum learning [82, 83, 84] may offer a solution. However, it typically focuses on selecting whole image samples, whereas MIM pre-training masks large portions of image pixels, emphasizing patch-level selection. Therefore, it’s essential to distinguish between “easy” and “hard” patches in MIM pre-training.
To this end, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection. In this module, depicted in Fig. 3, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set. This simple yet effective strategy ensures that semantically rich tokens are selected. Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens. This process is outlined in the Appendix. Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values. Recalling our partial reconstruction approach in Sec. 3.2.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets. Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process. Conversely, if a token’s feature distribution deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn. Details of the process are presented in the Appendix.
4 Experiment
In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments. We also conducted further experiments to test the scalability of the pipeline in the ablation experiments. Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks.
Pre-training Setup. Our pre-training experiment setup largely follows MAE [10]. The masking ratio is 85%, and the reconstruction ratio is 25%. For patch selection, the adjustment threshold is set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
4.1 Ablation Study
To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-4M dataset generated by the pipeline and the accelerated training method, SelectiveMAE. Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes.
4.1.1 Superiority of OpticalRS-4M
Table 1: Comparison of downstream task performance at matched data throughput between MillionAID and OpticalRS-4M, assessed using MAE [10] pre-training and ViT Base models [39].
Dataset Method Images Number Epoch Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MillionAID MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
OpticalRS-4M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
OpticalRS-4M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
Dataset Diversity. In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance. We define dataset diversity as its ability to improve various types of these tasks. While datasets like MillionAID focus on classification, our OpticalRS-4M dataset is designed to extend the benefit to detection and segmentation tasks. We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput. For instance, MillionAID, with 1 million samples, is trained using the basic MAE method for 800 epochs as a benchmark. To showcase the diversity of OpticalRS-4M, containing 4 million samples, we train it with MAE for just 200 epochs. Additionally, we validate OpticalRS-4M’s diversity by progressively reducing its training data — down to 3 million, 2 million, or even 1 million samples — while keeping data throughput constant. As shown in Table 1, our experiments demonstrate that even when randomly sampling 1 million images from OpticalRS-4M, it outperforms MillionAID in downstream tasks, particularly in object detection and segmentation. This consistent superior performance, even with fewer images, highlights OpticalRS-4M’s greater diversity and effectiveness, emphasizing the benefits of using larger and more varied datasets for unsupervised pre-training.
The results in Table 1 indicate that the optimal outcomes at equivalent data throughput are achieved with 3 million samples rather than 4 million. In most cases, the model pre-trained on 3 million samples outperforms across various downstream tasks. Notably, on the DIOR and LoveDA datasets, the best results were obtained with 1 million samples trained for 800 epochs. However, as shown in Table 2, when we extended the training on 4 million samples to more epochs, the model achieved the highest average performance across most downstream tasks. This suggests that training on 4 million samples for only 200 epochs was insufficient for the MIM method to fully leverage the large dataset, resulting in inferior generalization in downstream tasks.
Table 2: Performance on downstream tasks with extended training on OpticalRS-4M.
Dataset Method Images Number Epoch Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-4M MAE 4 million 100 95.12/97.54 91.26/93.85 - - - -
OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
OpticalRS-4M MAE 4 million 400 96.36/98.16 92.41/94.03 - - - -
OpticalRS-4M MAE 4 million 800 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
Extended training on OpticalRS-4M. To maximize the potential of OpticalRS-4M, we extended training using the MAE pre-training method and ViT-B, varying only the number of training epochs while keeping other settings consistent. The fine-tuning results, shown in Table 2, reveal that as the number of epochs increases, OpticalRS-4M shows greater advantages in downstream tasks. This indicates that extended training is necessary to fully harness OpticalRS-4M’s capabilities. However, the gains diminish with additional epochs, suggesting that merely increasing epochs may not suffice. To unlock the full potential of OpticalRS-4M, exploring more effective pre-training methods, such as the proposed SelectiveMAE, may be crucial.
Table 3: Comparison of downstream task performance using SelectiveMAE and MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
Method Dataset Epoch Total Training Sence Object Semantic
Classification Detection Segmentation
Hours AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MAE OpticalRS-4M 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE OpticalRS-4M 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
4.1.2 Effectiveness of SelectiveMAE
Performance Advantage. To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-4M using both MAE [10] and SelectiveMAE. The performance of ViT-B on various downstream tasks after pre-training is presented in Table 3. The results indicate that after pre-training on OpticalRS-4M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed. This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-4M.
Table 4: The accelerating performance of SelectiveMAE for different backbones compared to MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
Model Backbone Params Total Training Hours
MAE ViT-B 86M 101
SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×)
MAE ViT-L 307M 113
SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×)
Pre-training Efficiency. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the pretraining time with different backbones on the MillionAID dataset, as shown in Table 4. It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger sized models.
4.1.3 Scalability of the Pipeline
Table 5: Training cost (measured using 8 NVIDIA A100 GPUs) of SelectiveMAE and MAE [10] on OpticalRS-4M with different data volumes.
Method Images Number Epoch Total Training Hours
MAE 1 million 800 54
SelectiveMAE 1 million 800 25 (-29)
MAE 4 million 800 177
SelectiveMAE 4 million 800 86 (-81)
To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each. The results, presented in Table 5, show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours. When the dataset was expanded to 4 million samples, the time savings increased to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
4.2 Main Results
4.2.1 Fine-tuning Performance on Downstream Tasks
We further evaluated SelectiveMAE with the OpticalRS-4M dataset across three key downstream tasks: scene classification, object detection, and semantic segmentation. In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[54], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models. These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning-based approaches [56]. Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE). More details can be found in the Appendix.
Scene Classification. We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders. We leverage two scene classification datasets: AID [85] and RESISC-45 [86]. Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code. Evaluation is based on overall accuracy (OA). Table 6 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline. Besides, it outperforms RVSA and others when scaled to ViT-L. This demonstrates that, with our SelectiveMAE training method on the OpticalRS-4M dataset, the model can learn strong feature representations and scale efficiently.
Table 6: Results for scene classification, object detection, and semantic segmentation. “TR" represents the ratio of training data to the entire dataset.
Model Publication Backbone Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
SeCo [22] ICCV’21 ResNet-50 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [24] ICCV’21 ResNet-50 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [28] JSTARS’23 ResNet-50 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [21] CVPR’23 ResNet-50 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [30] NIPS’22 ViT-L 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [33] ICCV’23 ViT-L 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [54] GRSM’23 ViT-S 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [37] TGRS’22 Swin-B 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [56] ICCV’23 Swin-B 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [25] ICCV’23 Swin-B 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [18] TGRS’23 ViT-B+RVSA 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
SelectiveMAE - ViT-B 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
SelectiveMAE - ViT-L 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [87] and its improved variant DIOR-R for oriented object detection [88], both comprising RGB images. Following the methodologies of previous works [18, 37], we maintained consistent experimental setups, employing Faster-RCNN [89] and Oriented-RCNN [90] as detectors for each dataset. Detailed experimental configurations are available in the accompanying code. The results are summarized in Table 6. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [51] and SpaceNetv1 [52]. Our implementation closely follows [18], utilizing UperNet [91] as the segmentation framework, with mean Intersection over Union (mIOU) as the evaluation metric. Table 6 demonstrates the clear superiority of SelectiveMAE over its competitors in pixel-level semantic segmentation. The selection and reconstruction strategies employed by SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. Built upon it, we first curated OpticalRS-4M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-4M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-4M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that only 40% RS image patches are needed to train a comparable MIM model. Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE achieves over 2×2$\times$2 × speedup, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.