### v4

# Scaling Efficient Masked Image Modeling on Large Remote Sensing Dataset

Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-13M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-13M comprises 13 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments demonstrate that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models.
1 Introduction
Figure 1: Advantages of SelectiveMAE on finetuning performance and pre-training efficiency. The accuracy is evaluated on the DIOR detection dataset [41], while the training efficiency is measured on NVIDIA A100 GPUs, with different versions of ViT [19].
Over the past decade, advancements in remote sensing (RS) technology and data acquisition have improved applications in ecosystem monitoring [54], natural disaster management [6], among others [32, 74]. These applications rely on essential capabilities such as scene classification [30, 14], object detection [43], change detection [85], and semantic segmentation [93]. However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
Figure 2: Advantages of OpticalRS-13M in data volumn and diversity. (a) The volumn of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets. (b) OpticalRS-13M includes significantly more sample categories than others, spanning both object and event types. (c) OpticalRS-13M provides more abundant feature representations compared to other large-scale RS datasets. Here, 1,000 samples are separately selected from different datasets, and features are extracted using a pre-trained ViT-B.
Thanks to advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [28, 89], the pre-training of visual foundation models has seen remarkable improvements [90, 94, 91, 95, 73, 77]. Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [70]. However, two challenges persist in the development of RSFMs. (i) Compared to the ImageNet-21k [58] dataset, previous RS datasets [48, 50, 49, 62] contain fewer samples (approximately 1 million vs. 14 million), which impedes the sufficient MIM training of large backbones. (ii) These datasets, which primarily focus on global scene semantics [48, 50, 49], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks. This limitation restricts the generalization of the learned representations.
To address these challenges, taking an example of visible light RS imagery, we propose a new pipeline to collect, create, and efficiently process a large RS dataset. Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [48]. Nevertheless, issues such as inconsistent data sources, excessively large image sizes, and redundant pixels still exist. Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality. Correspondingly, we obtain a large-scale RS dataset called OpticalRS-13M comprising 13 million images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. OpticalRS-13M exceeds previous RS datasets [1, 50, 51, 33, 71, 65, 10, 17, 92, 35, 42, 56, 72, 64, 53, 70, 63, 52], being at least four times larger (Fig. 2 left). Moreover, OpticalRS-13M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 2 right).
Despite substantial efforts in training RSFM using MIM methods [51, 18, 17, 56], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [19] backbone on 8 Nvidia A100 GPUs [70]. This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-13M. In natural scene analysis, this issue has led to numerous studies [25, 75, 40, 11, 23, 20] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [25, 75]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [40, 11, 23], speeding up feature extraction.
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [70, 65]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches (e.g., $\leq$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens (e.g., $\leq$50%) with higher HOG values for feature encoding (e.g., $\leq$15%) and pixel reconstruction (e.g., $\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence (Fig. 1). Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
In summary, our main contributions are as follows:
1) We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
2) Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
3) We introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training. It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
Experiment results suggest the effectiveness and scalability of the proposed pipeline. OpticalRS-13M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [28].
2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [62, 84], object detection [16, 21, 38], and segmentation [2, 76, 68]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. The SEN12MS dataset [59] consists of 180,662 triplets, each containing synthetic aperture radar (SAR) and multispectral Sentinel-2 image patches, and MODIS land cover maps. While SSL4EO-S12 [79] , SSL4EO-L [61] and SatlasPretrain [5] contain millions of images, the major data type in these datasets are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [80] and RS5M [96], but they primarily focus on multimodal tasks. Currently, there are also some datasets that focus solely on visible light RS images. MillionAID [48], SeCo [49] and CACo [50] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, with the proposed pipeline, we introduce the OpticalRS-13M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [9], MIM has been developed for visual pre-training [12, 89, 28, 66, 37]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [29, 4, 82, 97, 47], masking strategies [60, 39], and reconstruction methods [67, 83, 45, 36]. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, approaches such as asymmetric encoder-decoder strategies [25, 75], reducing input patches [40, 23], or the Difficulty-Flatten Loss [11] have been proposed. Additionally, CrossMAE [20] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. Considering these issues, we introduce an efficient MIM method named SelectiveMAE that significantly speeds up pre-training, enhancing the practicality of developing RS foundation models based on large-scale datasets.
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [69]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [1, 50, 49, 81], recent advancements have primarily centered around generative self-supervised methods, especially in MIM. Specifically, many studies aim to improve generative self-supervised algorithms by leveraging general image knowledge [51, 33, 65], scaling up parameter sizes [10, 18], integrating spatio-temporal information [17, 92, 42], encompassing geometric attributes [70, 46], handling multi-sensor data [34, 24, 27, 8, 31], and employing multi-scale concepts [56, 64, 53]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
3 Method
The proposed pipeline contains two critical components: dataset generation and efficient pre-training, which will be introduced in detail in the following text.
3.1 Dataset Curation
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this, we curate a large-scale RS dataset, OpticalRS-13M, through extensive data collection and preprocessing.
Data Collection We began by reviewing publicly available RS datasets from the past decade. Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [48]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity. Intra-class Diversity: High within-class diversity ensures comprehensive representation of real-world target distributions. We include data from different sources of the same scene or target, incorporating change detection data to capture object variations by geographic location and imaging time. Inter-class Diversity: To enable self-supervised pretraining to distinguish categories effectively, we included as many fine-grained categories as possible, especially those with high semantic overlap, helping the model learn robust feature representations. 2) Richness: Beyond diversity, richness is essential. We ensure varied content features and large sample sizes by collecting images under multiple conditions (e.g., weather, seasons, lighting), introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion. Sampling across imaging conditions provides a comprehensive real-world representation, enhancing model’s representation and generalization capabilities. 3) Scalability: Scalability is crucial given the evolving applications of remote sensing images. We continuously expand the dataset based on scene and target categories across classification, detection, and segmentation tasks. In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios.
Data Preprocessing Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained. To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
1) We focused exclusively on visible light images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training. 2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices. 3) We filtered out images with fewer than 64 or more than 1024 pixels. 4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [55], and then a refined phase involving manual review. This ensured that only highly similar images were excluded. As a result of these steps, we created the OpticalRS-13M dataset, which contains approximately 13 million high-quality RS images, significantly larger than previous representative RS datasets, as illustrated in Fig. 2 left. In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality. We’ve screened and processed most small and medium-scale datasets from the past decade in this field.
Dataset Diversity To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig. 2 right. The analysis reveals that OpticalRS-13M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical remote sensing datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and so on. This expanded information enhances the versatility of OpticalRS-13M for a wider range of downstream tasks.
3.2 Efficient Pre-training
After obtaining a large-scale RS dataset, the next step is conduct the self-supervised pre-training. However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [28]. To address this issue, our pipeline introduces an efficient MIM method, SelectiveMAE. In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
3.2.1 Preliminaries of MAE
Figure 3: Overview of SelectiveMAE. It inputs fewer visible patches and reconstructs only partial patches to accelerate training efficiency.
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder. The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
However, applying MAE for self-supervised pre-training results in considerable training costs in terms of time complexity, especially when working with large RS datasets. To address this, we observe that RS optical images often contain redundant information, which could be omitted during training to improve pre-training efficiency. Specifically, we address two issues: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
3.2.2 Partial Reconstruction
For question 1, previous research [20] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches. In fact, reconstructing just 50% or even 25% of the patches can achieve similar performance and speed up training. However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones. Using only a random subset for reconstruction even degrades performance.
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 3.2.1. The remaining patches serve as reconstruction targets for the decoder.
Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
[…]
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [20]. Experimental results in supplementary material show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
3.2.3 Progressive Semantic Token Selection
Figure 4: Effectiveness of PSTS. Left: Using only 40% of patches for encoding and reconstruction often leads to gradient explosions. Right: The training loss after adopting PSTS.
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 3.2.2. However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig. 4 left. The figure shows that using an extremely low portion of patches for encoding and reconstruction caused unstable training due to gradient explosions.
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Inspired by the curriculum learning  [78, 7, 26], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig. 3. In this module, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. We define the initial set of selected token as:
[…]
We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set. This simple yet effective strategy ensures that semantically rich tokens are selected.
After token initialization, we incrementally increase the number of tokens to guide training from easier to more challenging examples, while maintaining the final masking ratio for selected tokens, as outlined in Algorithm 1. Given the high HOG feature values of the initial token sets, nearby tokens selected by PSTS also exhibit high HOG values. In partial reconstruction method in Section 3.2.2, we filter unselected tokens to retain those with high HOG values as reconstruction targets. Thus, selecting nearby tokens brings encoding tokens and reconstruction targets closer, streamlining training. In contrast, if a token’s feature distribution differs significantly from others, we consider these more challenging patches to learn.
Algorithm 1 Progressive Semantic Token Selection
1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
3:for $t\leftarrow 1$ to $T$ do
4:    Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
5:    $s\leftarrow\frac{1}{2}(1-m)$
6:    Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
7:    Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
[…]
9:    Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
10:end for
Specifically, we select tokens from $S^{U}$ based on $S^{I}$. First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer. We use Cosine Distance to measure the dissimilarity between the tokens in these two sets:
$\small\mathcal{D}(S^{U},S^{I})=\mathbf{1}-\cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-0.96873pt\left|\mathbb{S^{U}}\right|\kern-0.96873pt\right|\cdot\left|\kern-0.96873pt\left|\mathbb{S^{I}}\right|\kern-0.96873pt\right|},$ (3)
where $\mathbf{1}$ is an all-one matrix. $\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$. Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
[…]
where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs. Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
[…]
$\small{S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$. The operations in Eq. (5) and Eq. (6) are performed in each training epoch. This process is summarized in Algorithm 1. Fig. 5 further provides an example to illustrate which patches should be selected at different stages, helping to facilitate readers’ understanding.
Figure 5: Illustration of the patch selection process in PSTS.
4 Experiment
Table 1: Performance comparison of using the MillionAID and OpticalRS-13M datasets separately with the same data throughput, evaluated by MAE [28] and ViT-B [19].
Dataset Method Image Number Epoch Scene Object Semantic
Classification Detection Segmentation
AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MillionAID [48] MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99 75.20 66.99 52.56 79.43
OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60 74.50 66.36 52.63 79.39
In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments. We also conducted further experiments to test the scalability of the pipeline in the ablation experiments. Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks. All experiments are performed by fine-tuning pre-trained models on downstream tasks, including RS scene classification, object detection, and semantic segmentation. The pre-training and fine-tuning settings of our methods can be found in the supplementary material.
Table 2: The performance comparison between SelectiveMAE and the baseline method MAE [28] with the ViT-B backbone. The training time is measured on NVIDIA A100 GPUs. To accelerate validation, we select 4 million images from OpticalRS-13M for pre-training.
Method Image Number Epoch Total Training Scene Object Semantic
Classification Detection Segmentation
Time (h) AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MAE 4 million 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE 4 million 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
4.1 Ablation Study
To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated training method, SelectiveMAE. Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes. More ablation experiment results related to model design are presented in the supplementary material.
Dataset Diversity of OpticalRS-13M. In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance. Here, We define dataset diversity as the ability to improve performance across various types of tasks. While datasets like MillionAID [48] focus on classification, our OpticalRS-13M dataset is designed to extend the benefit to detection and segmentation tasks. We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput. As shown in Table 1, even with random sampling of 1 million images from OpticalRS-13M, it outperforms MillionAID on downstream tasks, particularly in object detection and segmentation. The results in Table 1 also indicate that the optimal outcomes at equivalent data throughput are achieved by different settings other than using all 13 million images. This suggests that training on a large number of samples with few epochs was insufficient for the MIM method to fully leverage the large dataset, leading to inferior generalization on downstream tasks. Experimental results in the supplementary material show consistent performance improvements across downstream tasks with more training epochs.
Table 3: The efficiency comparison between SelectiveMAE and the baseline method MAE [28] across different backbones on the MillionAID dataset [48] with 1600 epochs. The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
Model Backbone Params Total Training Time (h) GPU Memory (MB)
MAE ViT-B 86M 101 17628
SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×) 9570 (1.8×1.8$\times$1.8 ×)
MAE ViT-L 307M 113 30530
SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×) 18790 (1.6×1.6$\times$1.6 ×)
Performance Advantage of SelectiveMAE. To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-13M using both MAE [28] and SelectiveMAE. The performance of ViT-B on various downstream tasks after pre-training is presented in Table 2. The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed. This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
Efficiency Advantage of SelectiveMAE. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pretraining of different backbones on the MillionAID dataset, as shown in Table 3 (the corresponding accuracies are presented in Table 5). It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models. Additionally, it is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches.
Scalability of the Pipeline To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each. The results in Table 4 show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours. When the dataset was expanded to 4 million samples, the time savings increased to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
Table 4: The training cost of SelectiveMAE and the baseline method MAE [28] on OpticalRS-13M with different data volumes. All experiments are conducted on NVIDIA A100 GPUs.
Method Images Number Epoch Total Training Time (h)
MAE 1 million 800 54
SelectiveMAE 1 million 800 25 (-29)
MAE 4 million 800 177
SelectiveMAE 4 million 800 86 (-81)
Table 5: Performance comparison results of different models. “TR” represents the ratio of training data to the entire dataset. The first and second scores are marked by bold and blue, respectively. * means pre-training on 4 million images sampled from OpticalRS-13M.
Model Backbone Params (M) Scene Object Semantic
Classification Detection Segmentation
AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
SeCo [49] ResNet-50 26 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [1] ResNet-50 26 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [65] ResNet-50 26 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [50] ResNet-50 26 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [17] ViT-L 307 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [56] ViT-L 307 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [79] ViT-S 22 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [63] Swin-B 88 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [5] Swin-B 88 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [51] Swin-B 88 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [70] ViT-B+RVSA 86 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
LSKNet [44] LSKNet-S 14 97.05/98.22 94.27/95.83 - - 54.00 -
OREOLE [18] ViT-G 914 96.71/   -    -   /   - 77.40 71.31 54.00 -
SelectiveMAE* ViT-B 86 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
SelectiveMAE* ViT-L 307 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
SelectiveMAE ViT-B 86 96.64/98.30 92.53/94.50 75.10 67.69 52.68 79.44
SelectiveMAE ViT-L 307 97.21/98.48 94.16/95.90 78.70 71.75 53.92 79.48
4.2 Main Results
We compare SelectiveMAE to state-of-the-art RSFMs. In addition to benchmarking against MIM-based approaches such as SatMAE [17] and ScaleMAE [56], we also compare with contrastive learning-based methods, including GASSL [1] and SeCo [49], as well as advanced supervised learning models like SatLas [5]. Notably, none of these comparative methods were specifically designed for accelerated pre-training. Although their speed performance was not reported, we expect them to lag behind the baseline MAE.
Scene Classification. We first assess the pre-trained model’s performance on the scene classification task to evaluate its overall representational capability without additional decoders. We use two scene classification datasets: AID [86] and RESISC-45 [13], following training details and train-test splits as in [70, 56]. Table 5 shows that SelectiveMAE performs competitively against other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline. Additionally, it surpasses LSKNet and others when scaled to ViT-L, indicating that, with SelectiveMAE on the OpticalRS-13M dataset, the model effectively learns strong feature representations and scales efficiently.
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [41] and its improved variant DIOR-R for oriented object detection [15]. Following the methodologies of previous works [70, 63], we maintained consistent experimental setups, employing Faster-RCNN [57] and Oriented-RCNN [88] as detectors for each dataset. The results are summarized in Table 5. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [63]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE, which has close to 1B parameters, underscoring the excellent scalability of our method.
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [76] and SpaceNetv1 [68]. Our implementation follows [70], utilizing UperNet [87] as the segmentation framework. Table 5 demonstrates the clear superiority of SelectiveMAE over its competitors in semantic segmentation tasks. SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. Built upon it, we first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-13M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-13M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that using only 40% RS image patches is sufficient for training a comparable MIM model. Extensive experiments show that OpticalRS-13M significantly contributes to improving classification, detection, and segmentation performance, while SelectiveMAE achieves over a 2×2$\times$2 × speedup along with GPU memory savings, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.