### v5

# Harnessing Massive Satellite Imagery with Efficient Masked Image Modeling

Abstract
Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS). However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations. Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead. To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We curated a high-quality dataset named OpticalRS-13M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication. OpticalRS-13M comprises 13 million optical images covering various RS tasks, such as object detection and pixel segmentation. To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images. Extensive experiments show that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times. This highlights the effectiveness and scalability of our pipeline in developing RS foundational models. The dataset, source code, and trained models will be released at SelectiveMAE.
1 Introduction
Figure 1: Advantages of SelectiveMAE on finetuning performance and pre-training efficiency. The accuracy is evaluated on the DIOR detection dataset [1] with different versions of ViT [2]. The training efficiency is measured by Data Throughput/Minute, i.e., the processed images per minute, on 8 $\times$×NVIDIA A100 GPUs.
Over the past decade, advances in remote sensing (RS) technology and data acquisition have enhanced applications in ecosystem monitoring [3], natural disaster management [4], among others [5, 6]. These rely on essential capabilities such as scene classification [7, 8], object detection [9], change detection [10], and semantic segmentation [11]. However, each task typically demands substantial computational resources to train specialized models and learn task-specific feature representations.
Figure 2: Advantages of OpticalRS-13M in data volumn and diversity. (a) The volume of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets. (b) OpticalRS-13M includes significantly more sample categories than others, spanning both object and event types. y-axis denotes the data volume in (a-b). (c) OpticalRS-13M provides more diverse feature patterns than other large-scale RS datasets. Here, 1,000 samples are separately selected from different datasets, and their features from a pre-trained ViT-B are visualized using t-SNE [12].
Recent advances in self-supervised learning, particularly Masked Image Modeling (MIM) techniques [13, 14], have significantly improved the pre-training of visual foundation models [15, 16, 17, 18, 19, 20]. This progress has led to the emergence of remote sensing foundation models (RSFMs), which provide general feature representations and achieve state-of-the-art performance across various remote sensing tasks [21]. However, RSFMs still face two key challenges: (i) Compared to ImageNet-21k [22], existing RS datasets [23, 24, 25, 26] are significantly smaller (approximately 1 million vs. 14 million samples), limiting the effective MIM training of large models; and (ii) these datasets primarily capture global scene semantics [23, 24, 25] but lack the diversity and fine-grained details essential for downstream tasks, hindering the generalization of learned representations.
To address these challenges, taking an example of visible light RS imagery, we propose a new pipeline to collect, create, and efficiently process a large RS dataset. Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [23]. Nevertheless, issues such as inconsistent data sources, excessively large image sizes, and redundant pixels still exist. Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality. Correspondingly, we obtain a large-scale RS dataset called OpticalRS-13M comprising 13 million images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. OpticalRS-13M exceeds previous RS datasets [27, 24, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 21, 41, 42], being at least four times larger (Fig. 2 left). Moreover, OpticalRS-13M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 2 right).
Despite substantial efforts in training RSFM using MIM methods [28, 43, 33, 37], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [2] backbone on 8 Nvidia A100 GPUs [21]. This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-13M. In natural scene analysis, this issue has led to numerous studies [44, 45, 46, 47, 48, 49] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [44, 45]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [46, 47, 48], speeding up feature extraction.
Conventional MIM approaches, following the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [21, 31]. This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? 2) Is there a feasible way to encode fewer image patches (e.g., $\leq$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
Regarding the above issues, in the second part of the pipeline, we introduce an MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens (e.g., $\leq$50%) with higher HOG values for feature encoding (e.g., $\leq$15%) and pixel reconstruction (e.g., $\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion. To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase. In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence (Fig. 1). Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
In summary, our main contributions are as follows:
  1. 1.
    We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models. Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
  2. 2.
    For this pipeline, we introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training. It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
  3. 3.
    Experiment results demonstrate the effectiveness and scalability of the proposed pipeline. OpticalRS-13M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [13].
2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [26, 50], object detection [51, 52, 53], and segmentation [54, 55, 56]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works combine various sensor data to create extensive datasets. The SEN12MS dataset [57] consists of 180,662 triplets, each containing synthetic aperture radar (SAR) and multispectral Sentinel-2 image patches, and MODIS land cover maps. While SSL4EO-S12 [58] , SSL4EO-L [59] and SatlasPretrain [60] contain millions of images, the major data type in these datasets are multispectral and SAR, with only a small fraction being RGB. There are also some large-scale RS visual-language datasets such as Skyscript [61] and RS5M [62], but they primarily focus on multimodal tasks. Currently, there are also some datasets that focus solely on visible light RS images. MillionAID [23], SeCo [25] and CACo [24] provide nearly a million images of the same location over different times. However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, we introduce the OpticalRS-13M dataset, which is larger and more diverse. Using this dataset for pre-training, it significantly enhances performance across multiple downstream tasks.
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [63], MIM has been developed for visual pre-training [64, 14, 13, 65, 66]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [67, 68, 69, 70, 71], masking strategies [72, 73], and reconstruction methods [74, 75, 76, 77]. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, approaches such as asymmetric encoder-decoder strategies [44, 45], reducing input patches [46, 48], or the Difficulty-Flatten Loss [47] have been proposed. Additionally, CrossMAE [49] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. Considering these issues, we introduce an efficient MIM method named SelectiveMAE that significantly speeds up pre-training, enhancing the practicality of developing RS foundation models based on large-scale datasets.
Remote Sensing Foundation Models. Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [78]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [27, 24, 25, 79, 80], recent advancements have primarily centered around generative self-supervised methods, especially in MIM. Specifically, many studies aim to improve generative self-supervised algorithms by leveraging general image knowledge [28, 29, 31], scaling up parameter sizes [32, 43], integrating spatio-temporal information [33, 34, 36], encompassing geometric attributes [21, 81], handling multi-sensor data [82, 83, 84, 85, 86], and employing multi-scale concepts [37, 39, 40]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RSFMs, enhancing the practicality of MIM pre-training on large-scale datasets.
Figure 3: The comparison between OpticalRS-13M and existing RS large-scale datasets in terms of data volume and image type.
3 Method
The proposed pipeline contains two critical components: dataset generation and efficient pre-training, which will be introduced in detail in the following text.
3.1 Dataset Curation
Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets. To overcome this issue, we curate a large-scale RS dataset with diverse coverage scenarios, named OpticalRS-13M. Detailed data collection and preprocessing are presented in the supplementary material.
Figure 4: Overview of SelectiveMAE. It inputs fewer visible patches and reconstructs only partial patches to accelerate training efficiency.
Data Volume. OpticalRS-13M comprises 13 million high-resolution visible light RS images, establishing a significant scale advantage over existing datasets, as illustrated in Figure 3. OpticalRS-13M contains 13,203,698 images with a total of 2,630,362,174,503 pixels, averaging 199,214 pixels per image. We also offer a high-quality, lightweight version: OpticalRS-4M. OpticalRS-4M contains 3,920,829 images with a total of 688,459,799,359 pixels, averaging 175,590 pixels per image. While large-scale datasets such as SSL4EO-S12 [58] and SatLasPretrain [60] incorporate diverse data formats, e.g., including multispectral imagery, OpticalRS-13M is uniquely focused on visible light data. The dataset is carefully curated from multi-sensor acquisitions, leveraging imagery from high-resolution satellite platforms such as WorldView [87], QuickBird [88], GeoEye [89], and others. This specialization positions OpticalRS-13M as the most extensive visible light RS image repository to date, offering unparalleled utility for training vision foundation models for RS applications.
Dataset Diversity. To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig. 2 (b) and (c). The analysis reveals that OpticalRS-13M comprises 12 main categories, each containing numerous subclasses. Unlike existing self-supervised optical RS datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category. This category includes labels such as “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and so on. This expanded information enhances the versatility of OpticalRS-13M for a wider range of downstream tasks.
3.2 Efficient Pre-training
After obtaining a large-scale RS dataset, the next step is conduct the self-supervised pre-training. However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [13]. To address this issue, our pipeline introduces an efficient MIM method, SelectiveMAE. In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
3.2.1 Preliminaries of MAE
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder. The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error between the reconstructed targets and original patches.
However, applying MAE for self-supervised pre-training results in considerable training costs in terms of time complexity, especially when working with large RS datasets. To address this, we observe that RS optical images often contain redundant information, which could be omitted during training to improve pre-training efficiency. Specifically, we address two issues: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
3.2.2 Partial Reconstruction
For question 1, previous research [49] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches. In fact, reconstructing just 50% or even 25% of the patches can achieve similar performance and speed up training. However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones. Using only a random subset for reconstruction even degrades performance.
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 3.2.1. The remaining patches serve as reconstruction targets for the decoder.
Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
[…]
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [49]. Experimental results in supplementary material show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
3.2.3 Progressive Semantic Token Selection
Figure 5: Effectiveness of PSTS. Left: Using only 40% of patches for encoding and reconstruction often leads to gradient explosions. Right: The training loss after adopting PSTS.
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 3.2.2. However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig. 5 left. The figure shows that using an extremely low portion of patches for encoding and reconstruction caused unstable training due to gradient explosions.
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Inspired by curriculum learning  [90, 91, 92], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig. 4. In this module, we begin by selecting a limited number of patches and then select additional patches based on the similarity measure in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. $m$ is the mask ratio. We define the initial set of selected token as:
[…]
We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set. This simple yet effective strategy ensures that semantically rich tokens are selected.
After token initialization, we incrementally increase the number of tokens to guide training from easier to more challenging examples, while maintaining the final masking ratio for selected tokens, as outlined in Algorithm 1. Given the high HOG feature values of the initial token sets, nearby tokens selected by PSTS also exhibit high HOG values. In partial reconstruction method in Section 3.2.2, we filter unselected tokens to retain those with high HOG values as reconstruction targets. Thus, selecting nearby tokens brings encoding tokens and reconstruction targets closer, streamlining training. In contrast, if a token’s feature distribution differs significantly from others, we consider these more challenging patches to learn.
Figure 6: Illustration of the patch selection process in PSTS.
Algorithm 1 Progressive Semantic Token Selection
1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
3:for $t\leftarrow 1$ to $T$ do
4:    Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
5:    $s\leftarrow\frac{1}{2}(1-m)$
6:    Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
7:    Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
[…]
9:    Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
10:end for
Specifically, we select tokens from $S^{U}$ based on $S^{I}$. First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer. We use Cosine Distance to measure the dissimilarity between the tokens in these two sets:
$\mathcal{D}(\mathbf{S}^{U},\mathbf{S}^{I})=1-\cos\left\langle\mathbf{S}^{U},\mathbf{S}^{I}\right\rangle=1-\frac{\mathbf{S}^{U}(\mathbf{S}^{I})^{\mathrm{T}}}{\lVert\mathbf{S}^{U}\rVert\cdot\lVert\mathbf{S}^{I}\rVert}$ (3)
where $\mathbf{1}$ is an all-one matrix. $\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$. Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
[…]
where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs. Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
[…]
${S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$. The operations in Eq. (5) and Eq. (6) are performed in each training epoch. This process is summarized in Algorithm 1. Fig. 6 further provides an example to illustrate which patches should be selected at different stages, helping to facilitate readers’ understanding.
4 Experiment
In this section, we performed comprehensive comparative experiments to evaluate the performance of the proposed method on various downstream tasks, including RS scene classification, object detection, and semantic segmentation. Then, we validated the effectiveness of the OpticalRS-13M dataset and the SelectiveMAE method through ablation experiments. We also conducted further experiments to test the scalability of the pipeline. All experiments are performed by fine-tuning. The pre-training and fine-tuning settings of our methods can be found in the supplementary material.
Table 1: Performance comparison results of different models. “TR” represents the ratio of training data to the entire dataset. The first and second scores are marked by bold and blue, respectively. $\dagger$ means pre-training on 4 million images sampled from OpticalRS-13, with the epoch is set to 800. The overall accuracy is adoped as the metric for the scene classification task.
Model Backbone Params (M) Data Scene Object Semantic
Throughput Classification Detection Segmentation
/Minute AID [93] RESISC-45 [94] DIOR [1] DIOR-R [95] LoveDA [55] SpaceNetv1 [56]
TR=20%/50% TR=10%/20% mAP50 mAP50 mIoU mF1
SeCo [25] ResNet-50 [96] 26 - 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [27] ResNet-50 [96] 26 - 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [31] ResNet-50 [96] 26 - 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [24] ResNet-50 [96] 26 - 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [33] ViT-L [2] 307 205k 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [37] ViT-L [2] 307 206k 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [58] ViT-S [2] 22 - 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [41] Swin-B [97] 88 - 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [60] Swin-B [97] 88 243k 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [28] Swin-B [97] 88 - 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [21] ViT-B+RVSA [21] 86 - 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
OREOLE [43] ViT-G [43] 914 - 96.71/   -    -   /   - 77.40 71.31 54.00 -
MAE [13$\dagger$ ViT-B [2] 86 264k 96.58/98.02 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE $\dagger$ ViT-B [2] 86 556k 96.90/98.12 93.35/94.58 75.70 67.78 53.05 79.50
[…]
SelectiveMAE ViT-B [2] 86 556k 97.10/98.28 93.70/95.48 75.80 67.69 52.68 79.44
SelectiveMAE ViT-L [2] 307 533k 97.49/98.52 94.73/96.36 78.70 71.75 53.92 79.48
4.1 Main Results
We compare SelectiveMAE to state-of-the-art RSFMs. In addition to benchmarking against MIM-based approaches such as SatMAE [33] and ScaleMAE [37], we also compare with contrastive learning-based methods, including GASSL [27] and SeCo [25], as well as advanced supervised learning models like SatLas [60]. Notably, these comparative methods were specifically designed for increasing performance rather than accelerating pre-training. The Scene Classification and Detection comparison results are taken directly from SkySense [83], while those for Semantic Segmentation are exactly sourced from MTP [38]. Although their speed performance was not reported, we expect them to lag behind the baseline MAE. The performance of ViT-B and ViT-L on various downstream tasks after pre-training is presented in Table 1. The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE outperforms other RSFMs. Notably, it not only outperforms baseline (MAE) but also achieves this at 2.1 times the speed (556/264). This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
Scene Classification. We first assess the pre-trained model’s performance on the scene classification task to evaluate its overall representational capability without additional decoders. We use two scene classification datasets: AID [93] and RESISC-45 [94], following training details and train-test splits as in [21, 37]. Table 1 shows that SelectiveMAE performs competitively against other pre-training methods on both datasets. Additionally, when scaled to ViT-L, our model outperforms OREOLE [43] and other competitors, indicating the efficacy of SelectiveMAE pre-training on the OpticalRS-13M dataset, which enables strong feature representation learning while maintaining efficient scalability with increasing model size.
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [1] and its improved variant DIOR-R for oriented object detection [95]. Following the methodologies of previous works [21, 41], we maintained consistent experimental setups, employing Faster-RCNN [98] and Oriented-RCNN [99] as detectors for each dataset. The results are summarized in Table 1. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [41]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE [43], which has close to 1B parameters, underscoring the excellent scalability of our method.
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known RS datasets: LoveDA [55] and SpaceNetv1 [56]. Our implementation follows [21], utilizing UperNet [100] as the segmentation framework. Table 1 demonstrates the clear superiority of SelectiveMAE over its competitors in semantic segmentation tasks. SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
4.2 Ablation Study
To validate the efficacy of SelectiveMAE, we conducted ablation experiments about HOG Selecting, Similarity Measure and Selection Pattern. To further verify the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated performance of SelectiveMAE. Additionally, we compared the SelectiveMAE with the CrossMAE, a MAE-based method for natural images. More ablation experiment results are presented in the supplementary material.
Alternatives of HOG Selecting. To verify the efficacy of HOG in token selecting, we conduct two ablation experiments using MillionAID dataset [23] to replace HOG: (1) we used a pretrained Swin-B [97] model and applied k-nearest neighbor clustering to the extracted features. (2) We employed a lightweight feature extractor network based on Adamae [101]. Results in the Table 2 demonstrate HOG’s pre-training speed advantage. While using Swin-B slightly outperforms HOG, it is highly inefficient, and pre-training a foundation model like Swin requires massive data and computational resources. For RS images with abundant low-level features, HOG effectively selects tokens while ensuring substantially higher training speeds.
Table 2: Ablation experiment results of HOG selection through pre-training on the MillionAID dataset for 800 epochs.
Method Params Data Throughput AID RESISC-45
/ Minute TR=20%/50% TR=10%/20%
Adamae [101] 2.36M 498k 88.78/91.25 85.72/87.44
Swin-B 88M 356k 93.21/96.48 89.94/93.72
HOG - 556k 93.17/96.12 89.21/92.31
Ablation of Similarity Measure and Selection Pattern. To comprehensively assess the PSTS, we tested different Similarity Measures and the Selection Pattern. The results in the Table 3 reveal that different similarity measures had minimal impact on performance. However, changing the selection pattern led to a performance drop, confirming the effectiveness of the PSTS.
Table 3: Ablation experiment results of different similarity measures and selection patterns in PSTS through pre-training on the MillionAID dataset for 800 epochs.
Selection Pattern Similarity Measure AID RESISC-45
TR=20%/50% TR=10%/20%
near-far-random Euclidean distance 93.02/95.96 88.97/92.07
near-far-random Manhattan distance 92.92/95.89 89.17/91.92
far-near-random Cosine distance 90.12/92.78 85.81/88.80
near-far-random Cosine distance 93.17/96.12 89.21/92.31
Table 4: Performance comparison of using different pre-training settings on datasets, methods and epochs, where the ViT-B [2] is adopted as the network.
Dataset Method Image Number Epoch AID [93] RESISC-45 [94]
TR=20%/50% TR=10%/20%
Different Datasets
MillionAID [23] MAE 1 million 800 94.92/97.38 89.20/93.60
OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88
Equivalent Data Throughput
OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31
OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41
OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30
OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99
OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60
OpticalRS-13M SelectiveMAE 1 million 800 96.29/97.78 91.41/93.48
OpticalRS-13M SelectiveMAE 4 million 200 95.96/98.06 92.06/94.05
OpticalRS-13M SelectiveMAE 13 million 67 96.31/97.95 91.88/93.76
More Epoch
OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60
OpticalRS-13M MAE 13 million 100 96.34/98.13 91.79/93.85
OpticalRS-13M MAE 13 million 200 96.51/98.22 92.19/94.26
OpticalRS-13M MAE 13 million 800 97.10/98.28 93.70/95.48
Efficacy of OpticalRS-13M.
As illustrated in Table 4, when pre-training with a randomly sampled subset of 1 million images from OpticalRS-13M, the model achieves superior performance compared to MillionAID [23], underscoring the efficacy of OpticalRS-13M for representation learning. To further assess dataset diversity, we conducted experiments with equivalent data throughput during pre-training. The results in Table 4 reveal that optimal performance is attained by varying training configurations rather than utilizing the entire 13 million image corpus, regardless of the pre-training methodology. In our opinion, this observation highlights the inherent diversity of OpticalRS-13M, as training with fewer epochs proved insufficient for the MIM method to fully exploit the dataset’s potential, leading to model underfitting. To validate this hypothesis, we extended the training schedule, and the experimental results presented in the last part of Table 4 confirm our claim.
Table 5: The efficiency comparison between SelectiveMAE and the baseline method MAE [13] across different backbones on the part of OpticalRS-13M dataset [23] with 800 epochs. The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
Model Backbone Data Volume Training Time (h) GPU Memory (MB)
MAE ViT-B 1 million 51 17628
SelectiveMAE ViT-B 1 million 24 (2.1×2.1$\times$2.1 ×, -27) 9570 (1.8×1.8$\times$1.8 ×)
MAE ViT-L 1 million 57 30530
SelectiveMAE ViT-L 1 million 25 (2.3×2.3$\times$2.3 ×, -32) 18790 (1.6×1.6$\times$1.6 ×)
MAE ViT-B 4 million 177 17628
SelectiveMAE ViT-B 4 million 86 (2.1×2.1$\times$2.1 ×, -81) 9570 (1.8×1.8$\times$1.8 ×)
Efficiency Advantage of SelectiveMAE. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pre-training of different backbones on the part of OpticalRS-13M dataset, as shown in Table 5 (the corresponding accuracies are presented in Table 1). It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models. It is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches. To further evaluate the scalability of our pipeline, we organized datasets containing 4 million samples and conducted experiments. The results in Table 5 show that, when the dataset was expanded to 4 million samples, the time savings increased from 27 (51-24) to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
Comparison with CrossMAE [49]. CrossMAE, an MAE-based self-supervised pretraining method, is originally designed for natural images. Since SelectiveMAE adopts the same decoder as CrossMAE [49] during pre-training, and incorporates a HOG-based strategy tailored to RS images inspired by its partial reconstruction concept, as detailed in Section 3.2.2. In this section, we have also complemented related comparison experiments. Nevertheless, due to CrossMAE has not been pretrained in the RS field, we conducted a fair evaluation through pretraining it on the OpticalRS-13M dataset using the official code. Table 6 shows that SelectiveMAE outperforms CrossMAE in both accuracy and efficiency. By integrating RS-specific enhancements, PSTS and partial reconstruction, SelectiveMAE proves to be a highly effective self-supervised learning approach for RS.
Table 6: Performance comparison between our method and CrossMAE, where 400 million images of OpticalRS-13M are randomly sampled for pre-training 800 epochs.
Method Data Throughput / Minute AID RESISC-45
TR=20%/50% TR=10%/20%
CrossMAE(ViT-B) 475k 96.12/97.18 92.08/94.12
SelectiveMAE(ViT-B) 556k 96.78/98.12 93.35/94.58
5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-13M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-13M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that using only 40% RS image patches is sufficient for training a comparable MIM model. Extensive experiments show that OpticalRS-13M significantly contributes to improving classification, detection, and segmentation performance, while SelectiveMAE achieves over a 2×2$\times$2 × speedup along with GPU memory savings, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.